CORPORATE VALUATION

Dantzler Corporation is a fast-growing supplier of office products. Analysts project the following free cash flows (FCFs) during the next 3 years, after which FCF is expected to grow at a constant 6% rate. Dantzler's WACC is 12%.

Year 0 1 2 3
....... ....... ....... ....... ....... ....... ....... .......
....... ....... ....... ....... ....... ....... ....... ......
FCF ($ millions) - $8 $21 $45
  1. What is Dantzler's horizon, or continuing, value? (Hint: Find the value of all free cash flows beyond Year 3 discounted back to Year 3.) Round your answer to two decimal places. Enter your answer in millions. For example, an answer of $13,550,000 should be entered as 13.55.
    $ million
  2. What is the firm's value today? Round your answer to two decimal places. Enter your answer in millions. For example, an answer of $13,550,000 should be entered as 13.55. Do not round your intermediate calculations.
    $ million
  3. Suppose Dantzler has $57 million of debt and 32 million shares of stock outstanding. What is your estimate of the current price per share? Round your answer to two decimal places. Write out your answer completely. For example, 0.00025 million should be entered as 250.
    $

Please help!

Thank you!!
Answer #1

1.
=FCF3*(1+g)/(WACC-g)
=45*1.06/(12%-6%)
=795.00 million

2.
=-8/1.12+21/1.12^2+45/1.12^3+45/1.12^3*1.06/(12%-6%)
=607.4936224 million

3.
=(Enterprise Value-Debt)/Number of shares
=(607.4936224-57)/32
=17.2029257